OpenStudy (anonymous):

Find the limit of the function algebraically: lim x-->0 (x^2+3)/x^4

OpenStudy (anonymous):

@TheSmartOne @uri

TheSmartOne (thesmartone):

I'm not sure, I haven't learned this, maybe @Michele_Laino @texaschic101 @johnweldon1993 @rishavraj can help you :)

OpenStudy (anonymous):

could i factor and then cancel? x(x+3)/x^4=

OpenStudy (anonymous):

x^3(x+3) is this right so far

OpenStudy (johnweldon1993):

Well the original question is \[\large \lim_{x\rightarrow 0} \frac{x^2 + 3}{x^4}\] right?

OpenStudy (anonymous):

yes

OpenStudy (johnweldon1993):

Well I can tell you the top cannot be factored... you cannot take an X out because the 3 doesnt have an 'x' The bottom cannot be simplified in a way that would help anything here So if nothing can be simplified all we can do is plug 0 in for 'x' and see that \[\large \lim_{x \rightarrow 0} \frac{x^2 + 3}{x^4} = \infty\]

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so thats the answer?

OpenStudy (johnweldon1993):

Correct

OpenStudy (anonymous):

thanks so much!!!

OpenStudy (anonymous):

:')

OpenStudy (johnweldon1993):

Not a problem!